# Self-Data Distillation for Recovering Quality in Pruned Large Language Models

### v2

@@ 3 Ablation Studies on Layer-Pruning and Self-Data Distillation
The results, presented in Figure 3, show that self-data distillation offers significant quality improvements over SFT and one-shot pruned models, particularly with larger datasets like the 50k-sample OpenMathInstruct. Self-data distilled models achieved 5-10% higher accuracy across all pruning block sizes, demonstrating the effectiveness of self-data distillation in recovering a substantial portion of the model’s quality post-pruning. Our ablations revealed a strong correlation between dataset size and quality recovery, with larger datasets consistently outperforming smaller ones. The improvements being most pronounced with medium to large block sizes, where self-data distillation notably enhanced generalization, especially in reasoning tasks. These findings highlight the importance of dataset scale and the self-data distillation process, with the 50k-sample OpenMathInstruct dataset delivering the best overall results, and will be the focus of subsequent experiments.
Mitigating Catastrophic Forgetting
@@
The right plot in Figure 2 clearly illustrates the advantages of self-data distilled fine-tuning (Self-Data FT) over SFT in mitigating catastrophic forgetting after pruning. We study the similarities between Llama3.1-8B Instruct models fine-tuned on a supervised 50k-sample OpenMathInstruct dataset and a self-data distilled version of the same dataset. The plot compares the sentence embeddings of the two models’ generated responses on the GSM8k task to the baseline unpruned Llama3.1-8B Instruct model. Self-Data FT maintains a narrower distribution of embedding similarities, with a higher mean score of 0.92, indicating better preservation of the original model’s learned representations. In contrast, SFT exhibits a wider spread of similarity scores and a lower mean of 0.83, reflecting a more significant distribution shift and a greater risk of quality degradation. The tighter distribution in Self-Data FT highlights its ability to retain model quality across tasks, while the wider distribution in SFT suggests a greater distribution shift, leading to higher risk of catastrophic forgetting (see Appendix C). This highlights Self-Data FT as a more effective method for mitigating model quality degradation post-pruning.
@@ 4 Empirical Results
We evaluated the quality of Llama3.1-8B Instruct models pruned at various block sizes under three fine-tuning strategies: no fine-tuning (No FT), supervised fine-tuning (SFT), and our proposed self-data distillation. As shown in Table 1, pruned models without fine-tuning experience substantial accuracy losses, particularly at larger block sizes (e.g., a 46.39% average score at block size 10), highlighting the critical need for post-pruning adaptation. SFT improves quality, with an average recovery of 81.66% at block size 6, but struggles on reasoning-heavy tasks like GSM8k and ARC-C. In contrast, self-data distillation significantly enhances quality recovery, achieving 91.24% at block size 6, with GSM8k accuracy reaching 66.64% (compared to 43.82% with SFT). Even at block size 10, self-data distillation maintains an 80.56% recovery, outperforming SFT’s 68.56%. These findings establish self-data distillation as the most effective method for preserving model quality post-pruning, particularly for reasoning-intensive tasks, making it essential for large-scale model compression.
Improving Self-Data Distillation with Model Merging
We extend self-data distillation by introducing model merging using Spherical Linear Interpolation (SLERP) (Shoemake, 1985). While self-data distillation has shown significant improvements in maintaining model quality post-pruning, we investigate whether merging models fine-tuned on different datasets can provide further gains. Specifically, we merge Llama3.1-8B Instruct models fine-tuned on OpenMathInstruct and Alpaca, as these datasets delivered the best results in our ablations in Section 3. SLERP enables smooth interpolation between model parameters, preserving the geometric properties of the parameter space (see Appendix D for details). Our results, as shown in Table 1, demonstrate that self-data distillation with model merging via SLERP yields the highest recovery in quality across all pruning block sizes. At block size 6, the merged model achieves a 93.30% recovery, compared to 91.24% for the OpenMathInstruct model alone. These findings suggest that SLERP-based model merging not only mitigates pruning-related quality loss but also improves generalization, particularly on tasks such as GSM8k and ARC-C.
@@ 5 Related Work
To address these challenges, several metrics have been developed to guide pruning decisions more effectively. For instance, Shortened Llama (Kim et al., 2024) demonstrated that depth pruning (removing layers) can be as effective as width pruning (removing units within layers), or even a combination of both. The Block Influence (BI) score (Men et al., 2024), applied in Llama-2 (Touvron et al., 2023b), measures block importance by evaluating changes in hidden state magnitudes. Additionally, the angular cosine similarity metric (Gromov et al., 2024) identifies layers with redundant activations, allowing for selective pruning in models such as Llama-2 and Mistral (Jiang et al., 2023). Gromov et al. (2024) also proposed a healing method using low-rank adapters (Hu et al., 2022) to recover lost quality. Despite these advancements, pruning LLMs still results in sharp accuracy degradation (Sun et al., 2024), and traditional recovery methods such as fine-tuning or re-pretraining are resource-intensive (Xia et al., 2024; Sreenivas et al., 2024). Our self-data distillation approach extends this by leveraging the unpruned model to generate a distilled dataset for fine-tuning the pruned model, ensuring semantic alignment and mitigating the quality degradation caused by pruning. While combining this with standard KD techniques could further improve generalization, we leave that exploration for future work.
@@
Knowledge distillation (KD) (Hinton et al., 2015) is a widely-used model compression technique where a smaller student model learns from a larger teacher model, enabling efficient model quality retention. In NLP, KD has been applied in various contexts to align student models with teacher outputs (Liang et al., 2021; Gu et al., 2024; Agarwal et al., 2024), hidden states (Jiao et al., 2020), and attention mechanisms (Wang et al., 2021). Recent work on Llama3.2 (Llama Team, 2024) extends this by using logits from larger Llama3.1 models (e.g., 8B, 70B) as token-level targets during pre-training, allowing the smaller models (e.g., 1B, 3B) to achieve superior quality compared to training from scratch (Llama Team, 2024). Supervised fine-tuning (SFT) has been widely employed in various self-distillation frameworks to train student models using sequences generated by teacher LLMs (Sun et al., 2023; Wang et al., 2023; Zelikman et al., 2022). Yang et al. (2024) investigated self-distillation as a way to alleviate distribution shifts, improving model quality during SFT while improving generalization across tasks. Our self-data distillation method builds on these techniques by leveraging the original unpruned model to generate a distilled dataset for fine-tuning the pruned model. This ensures semantic alignment and mitigates the quality degradation seen after pruning. Furthermore, while our approach can be combined with standard KD techniques to enhance generalization and recover quality while lowering computational costs, we leave the exploration of such combinations for future work.
@@
One of the major challenges of pruning and distillation techniques in LLMs is catastrophic forgetting, where a model loses its previously learned capabilities during fine-tuning (Kotha et al., 2024; Korbak et al., 2022). Regularization techniques such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) aim to alleviate this by controlling parameter updates, but are task-dependent and require careful tuning (Huang et al., 2021). Architecture-based methods, which allocate separate parameters for each task (Razdaibiedina et al., 2023), preserve task-specific knowledge but add complexity and overhead, reducing the overall efficiency of model compression. Replay-based techniques (Ostapenko et al., 2022; Rolnick et al., 2019; Sun et al., 2019) store data subsets from previous tasks for rehearsal, either through direct storage or synthesis via generative models. However, these methods demand substantial memory to store large datasets and are often impractical due to privacy concerns or lack of access to past data. Our self-data distilled fine-tuning approach avoids these challenges by aligning the fine-tuning dataset with the original model’s learned distribution, preserving knowledge across tasks without requiring new parameters or architectural changes. This method offers a robust solution for mitigating catastrophic forgetting while maintaining model quality after pruning.

### v3

@@ 3.2 Analysis on Self-Data Distilled Datasets
The results, presented in Figure 2, show that self-data distillation offers significant quality improvements over SFT and one-shot pruned models, particularly with larger datasets like the 50k-sample OpenMathInstruct. Self-data distilled models achieve 5-10% higher accuracy than models without any fine-tuning for pruning block sizes above four. Across all block sizes, they also outperform models fine-tuned with SFT, demonstrating the effectiveness of self-data distillation in recovering a substantial portion of the model’s quality post-pruning. Our ablations revealed a strong correlation between dataset size and quality recovery, with larger datasets consistently outperforming smaller ones. The improvements being most pronounced with medium to large block sizes, where self-data distillation notably enhanced generalization, especially in reasoning tasks. These findings highlight the importance of dataset scale and the self-data distillation process, with the 50k-sample OpenMathInstruct dataset delivering the best overall results, and will be the focus of subsequent experiments.
3.3 Mitigating Catastrophic Forgetting
@@
Figure 4 illustrates the advantages of self-data distilled fine-tuning (Self-Data FT) over SFT in mitigating catastrophic forgetting after pruning. We study the similarities between Llama3.1-8B Instruct models fine-tuned on a supervised 50k-sample OpenMathInstruct dataset and a self-data distilled version of the same dataset. The plots compares the sentence embeddings of the two models’ generated responses on the GSM8k task to the baseline unpruned Llama3.1-8B Instruct model at various prune block sizes. Self-Data FT maintains a narrower distribution of embedding similarities, with higher mean scores, indicating better preservation of the original model’s learned representations. In contrast, SFT yields a wider spread and reduced mean in similarity scores, indicative of a heavier-tailed distribution. This shift suggests a more pronounced distributio shift, thereby increasing the risk of quality degradation. The tighter distribution in Self-Data FT highlights its ability to retain model quality across tasks, while the wider distribution in SFT suggests a greater distribution shift, leading to higher risk of catastrophic forgetting (see Appendix C). This highlights Self-Data FT as a more effective method for mitigating model quality degradation post-pruning.
Table 3: Model quality results for pruned Llama3.1-8B Instruct models across various pruning block sizes. Average accuracy is reported on the OpenLLM Leaderboard v1, along with the recovery percentage relative to the baseline model. Self-data distillation with model merging (MM) consistently outperforms self-data distillation with mixed datasets.
Prune Block Size Fine-tuning Method Dataset Avg. Score Avg. Recovery Baseline No FT 69.40 - 4 No FT 64.06 92.31% Self-Data Distillation OpenMathInstruct + Alpaca 65.27 94.04% Self-Data Distillation + MM OpenMathInstruct + Alpaca 65.84 94.86% 6 No FT 51.82 74.67% Self-Data Distillation OpenMathInstruct + Alpaca 63.77 91.88% Self-Data Distillation + MM OpenMathInstruct + Alpaca 64.75 93.30% 8 No FT 48.93 70.50% Self-Data Distillation OpenMathInstruct + Alpaca 59.37 85.54% Self-Data Distillation + MM OpenMathInstruct + Alpaca 61.24 88.24% 10 No FT 46.39 66.83% Self-Data Distillation OpenMathInstruct + Alpaca 55.85 80.47% Self-Data Distillation + MM OpenMathInstruct + Alpaca 56.08 80.70%
Table 4: Model quality results for pruned Mistral-7B-v0.3 Instruct models across various pruning block sizes. Average accuracy is reported on the OpenLLM Leaderboard v1, along with the recovery percentage relative to the baseline model. Self-data distillation with model merging (MM) consistently outperforms self-data distillation with mixed datasets.
Prune Block Size Fine-tuning Method Dataset Avg. Score Avg. Recovery Baseline No FT 66.11 - 4 No FT 56.90 86.06% Self-Data Distillation OpenMathInstruct + Alpaca 59.04 89.30% Self-Data Distillation + MM OpenMathInstruct + Alpaca 59.57 90.10% 6 No FT 52.77 79.82% Self-Data Distillation OpenMathInstruct + Alpaca 56.26 85.09% Self-Data Distillation + MM OpenMathInstruct + Alpaca 56.41 85.32% 8 No FT 49.53 74.92% Self-Data Distillation OpenMathInstruct + Alpaca 52.78 79.83% Self-Data Distillation + MM OpenMathInstruct + Alpaca 54.14 81.89% 10 No FT 41.00 62.01% Self-Data Distillation OpenMathInstruct + Alpaca 48.86 73.90% Self-Data Distillation + MM OpenMathInstruct + Alpaca 50.04 75.69%
@@ 4 Empirical Results
We evaluated the quality of Llama3.1-8B Instruct and Mistral-7B-v0.3 Instruct models pruned at various block sizes under three fine-tuning strategies: no fine-tuning (No FT), supervised fine-tuning (SFT), and our proposed self-data distillation. As shown in Table 1, pruned models without fine-tuning experience substantial accuracy losses, particularly at larger block sizes (e.g., a 46.39% average score at block size 10), highlighting the critical need for post-pruning adaptation. SFT improves quality, with an average recovery of 81.66% at block size 6, but struggles on reasoning-heavy tasks like GSM8k and ARC-C. In contrast, self-data distillation significantly enhances quality recovery, achieving 91.24% at block size 6, with GSM8k accuracy reaching 66.64% (compared to 43.82% with SFT). Even at block size 10, self-data distillation maintains an 80.56% recovery, outperforming SFT’s 68.56%. Similarly, in Table 2, the Mistral-7B models exhibit improved quality recovery across all pruning block sizes. These findings establish self-data distillation as an effective method for preserving model quality post-pruning, particularly for reasoning-intensive tasks, making it essential for large-scale compression.
Table 5: Evaluation of speculative decoding performance using self-data distilled Llama3.1-8B Instruct draft models for the Llama3.1-70B Instruct target model. The draft models generate 6 speculative tokens plus 1 base token from the target model, with accepted lengths measured across all tasks from Spec-Bench. Results are shown for various pruning block sizes, where self-data distillation with MM consistently improves over the no fine-tuning (No FT) baseline, with higher average accepted lengths indicating better speculative capabilities and better alignment with the target model. At a prune block size of 10, we achieve 27.16% FLOP savings, and the average accepted length across tasks (final column) improves by 1.70.
[…]
Table 6: Evaluation of speculative decoding performance using self-data distilled Mistral-7B-v0.3 Instruct draft models for the Mistral Large 2 (123B) target model. Results on Spec-Bench are shown for various pruning block sizes, highlighting the effectiveness of self-data distillation paired with model merging (MM). Self-data distillation with MM consistently improves over the no fine-tuning baseline, with higher average accepted lengths indicating better speculative capabilities and better alignment with the target model. At a prune block size of 10, we achieve 30.10% FLOP savings, and the average accepted length across tasks (final column) improves by 1.62.
[…]
4.1 Self-Data Distillation with Model Merging
We extend self-data distillation by introducing model merging using Spherical Linear Interpolation (SLERP) (Shoemake, 1985). The model merging technique aims to synthesize complementary knowledge from models fine-tuned on distinct datasets, potentially leading to improved generalization and robustness after pruning. While various model merging techniques, such as TIES (Yadav et al., 2023), DARE-TIES (Yu et al., 2024), and Linear Merge (Nagarajan & Kolter, 2019), have been proposed, prior empirical studies have demonstrated SLERP to be particularly effective (Aakanksha et al., 2024), as it effectively balances the dual objectives of maintaining alignment and preserving overall model quality. SLERP enables smooth interpolation between two model parameters along the shortest path on a high-dimensional sphere, preserving the geometric properties of the parameter space (see Appendix D for details).
Building on the gains achieved by self-data distillation in recovering quality post-pruning, we investigate whether merging models fine-tuned on diverse datasets can provide further improvements. We perform model merging within each model family by merging Llama3.1-8B Instruct models fine-tuned on OpenMathInstruct and Alpaca, and similarly merging Mistral-7B models fine-tuned on the same datasets, as these configurations delivered the best results in our ablations in Section 3. To establish a robust baseline for comparison against model merging, we combined the OpenMathInstruct and Alpaca datasets, fine-tuning the pruned models on this interleaved dataset. The total number of training samples was kept consistent across both approaches to ensure a fair evaluation.
As presented in Table 3 for Llama3.1 models and Table 4 for Mistral-7B models, our results indicate that self-data distillation combined with model merging via SLERP achieves the highest recovery of model quality across all pruning block sizes. Specifically, at a pruning block size of 6, the merged Llama3.1 model demonstrates a 93.30% recovery of its original quality, surpassing the 91.88% recovery observed with models trained on a mixed dataset. Similarly, the Mistral-7B model achieves an 85.32% recovery at block size 6 through merging, slightly exceeding the 85.09% recovery obtained with the mixed dataset approach. These results highlight the consistency and effectiveness of self-data distillation with model merging in retaining model quality across varying pruning configurations. These findings suggest that SLERP-based model merging not only mitigates pruning-related quality loss but also improves generalization, particularly on tasks such as GSM8k and ARC-C.
4.2 Speculative Decoding and Self-Data Distillation
Speculative decoding is founded on two key insights into the inference dynamics of LLMs. First, generating tokens with a smaller draft model reduces computational cost by offloading initial token prediction from the larger target model. Second, LLM inference is primarily constrained by memory bandwidth, with latency bottlenecks arising from parameter memory access rather than arithmetic operations (Patterson, 2004; Shazeer, 2019). By adapting principles from speculative execution, speculative decoding aims to prioritize the verification of pre-drafted tokens, thereby minimizing frequent memory operations and enhancing overall inference efficiency. Despite its potential, speculative decoding raises several challenges that require deeper exploration. Central among these is the need to design an effective drafting mechanism that optimally balances speculative accuracy with drafting efficiency (Xia et al., 2023; Li et al., 2024). The effectiveness of this approach depends on two critical factors: the accuracy of the draft model, measured by the average number of accepted tokens per decoding step, and the drafting latency itself (Stern et al., 2018; Xia et al., 2023). Achieving a trade-off between high speculative accuracy and low latency remains a significant challenge, as both elements are essential for maximizing the overall speedup.
While not the main focus of this work, speculative decoding is a natural application of our method. In this context, integrating self-data distillation with speculative decoding presents a viable solution to improve inference performance. Self-data distillation aligns the draft model more closely with the target model by generating a distilled dataset that preserves semantic consistency even after pruning. This integration not only enhances the draft model’s speculative capabilities but also mitigates quality degradation, allowing for more efficient and accurate token validation during inference. Consequently, the combination of speculative decoding and self-data distillation emerges as a promising approach for improving both inference speed and model quality retention across diverse tasks.
The draft-and-verify setup in speculative decoding involves two phases: drafting and verification. In the drafting phase, a smaller draft model $M_{p}$ speculates multiple future tokens. Given an input sequence $\{x_{1},\ldots,x_{n}\}$, it generates $K$ speculative tokens $\tilde{x}_{j}$ by sampling from probability distributions $p_{j}$, where $\{p_{1},\ldots,p_{K}\}=\text{Draft}(x_{\leq n},M_{p})$ and $\tilde{x}_{j}\sim p_{j}$ for $j=\{1,\ldots,K\}$. In the verification phase, the larger target model $M_{q}$ verifies these tokens by computing probabilities for the drafted tokens as $q_{j}=M_{q}(x\mid x_{\leq n},\tilde{x}_{<j})$ for $j=\{1,\ldots,K+1\}$. Each token $\tilde{x}_{j}$ is then checked against a criterion $\text{Verify}(\tilde{x}_{j},p_{j},q_{j})$, where in our experimental setup, verification is performed using $\tilde{x}_{j}=\arg\max q_{j}$. If a token fails verification, the first failing token $\tilde{x}_{c}$ is corrected via $\text{Correct}(p_{c},q_{c})$, specifically as $x_{n+c}\leftarrow\arg\max q_{c}$, and subsequent tokens are discarded. If all tokens pass verification, an additional token $x_{n+K+1}$ is sampled from the target model. This process repeats until the [EOS] token is generated or a maximum sequence length is reached.
Experimental Setup
Our speculative decoding experiments use pruned, self-data distilled models with different pruning block sizes $n\in\{4,6,8,10\}$. These models are evaluated on the Spec-Bench dataset (Xia et al., 2024a), a comprehensive benchmark that covers categories including multi-turn chat, translation, summarization, question answering (QA), mathematical reasoning, and retrieval-augmented generation (RAG). There are a total of 480 samples in Spec-Bench, where each category in Spec-Bench consists of 80 randomly selected instances drawn from six well-established datasets: MT-Bench (Zheng et al., 2024), WMT14 DE-EN (Nallapati et al., 2016), CNN/Daily Mail (Nallapati et al., 2016), Natural Questions (Kwiatkowski et al., 2019), GSM8K (Cobbe et al., 2021), and DPR (Karpukhin et al., 2020).
We report the overall average accepted token lengths to measure the effectiveness of speculative decoding. Our speculative decoding configuration generates $K=6$ speculative tokens $\{\tilde{x}_{1},\ldots,\tilde{x}_{6}\}$ using the draft model $M_{p}$, while the target model $M_{q}$ concurrently verifies these tokens and generates an additional token $x_{n+1}$. We explore two setups, 1)))) pruned self-data distilled Llama3.1 models speculating for the Llama3.1-70B Instruct (Dubey et al., 2024) target model, and 2)))) pruned self-data distilled Mistral-7B Instruct models speculating for the Mistral Large 2 (123B) (Mistral, 2024) target model.
Results
Our evaluations indicate that self-data distillation enhances speculative capabilities by improving alignment with the target model. Notably, with larger prune block sizes (e.g., 10), self-data distillation paired with model merging achieves higher average accepted lengths across all categories, demonstrating better generalization and more efficient speculative decoding. For instance, in Table 5 pruned Llama3.1 models at block size 10 (i.e., 5.85B resulting in 27.16% FLOPS savings) show an average accepted length improvement of 1.70 tokens. Similarly, in Table 6, we observe Mistral models at block size 10 (i.e., 5.06B resulting in 30.10% FLOPs savings) achieve a gain of 1.62 tokens. While the improvements are evident across general tasks, substantial gains are particularly observed in mathematical reasoning. Pruned Llama3.1 and Mistral-7B models at block size 10 demonstrate an average accepted length increase of 2.96 and 2.75 tokens, respectively, on mathematical reasoning benchmarks. These results show the effectiveness of self-data distillation in preserving complex reasoning capabilities, even under aggressive pruning. Such findings highlight the interplay between speculative decoding and self-data distillation, where the combination optimizes inference efficiency and model quality retention across a broad range of tasks, with notable robustness in challenging reasoning domains.
@@ 5 Related Work
To address these challenges, several metrics have been developed to guide pruning decisions more effectively. For instance, Shortened Llama (Kim et al., 2024) demonstrated that depth pruning (removing layers) can be as effective as width pruning (removing units within layers), or even a combination of both. The Block Influence (BI) score (Men et al., 2024), applied in Llama-2 (Touvron et al., 2023b), measures block importance by evaluating changes in hidden state magnitudes. Additionally, the angular cosine similarity metric (Gromov et al., 2024) identifies layers with redundant activations, allowing for selective pruning in models such as Llama-2 and Mistral (Jiang et al., 2023). Gromov et al. (2024) also proposed a healing method using low-rank adapters (Hu et al., 2022) to recover lost quality. Despite these advancements, pruning LLMs still results in sharp accuracy degradation (Sun et al., 2024), and traditional recovery methods such as fine-tuning or re-pretraining are resource-intensive (Xia et al., 2024b; Sreenivas et al., 2024). Our self-data distillation approach extends this by leveraging the unpruned model to generate a distilled dataset for fine-tuning the pruned model, improving semantic alignment and mitigating the quality degradation caused by pruning. While combining this with standard KD techniques could further improve generalization, we leave that for future work.
@@
Knowledge distillation (Hinton et al., 2015) is a widely-used model compression technique where a smaller student model learns from a larger teacher model, enabling efficient model quality retention. In NLP, KD has been applied in various contexts to align student models with teacher outputs (Liang et al., 2021; Gu et al., 2024; Agarwal et al., 2024), hidden states (Jiao et al., 2020), and attention mechanisms (Wang et al., 2021). Recent work on Llama3.2 (Llama Team, 2024) extends this by using logits from larger Llama3.1 models (e.g., 8B, 70B) as token-level targets during pre-training, allowing the smaller models (e.g., 1B, 3B) to achieve superior quality compared to training from scratch (Llama Team, 2024). Supervised fine-tuning (SFT) has been widely employed in various self-distillation frameworks to train student models using sequences generated by teacher LLMs (Sun et al., 2023; Wang et al., 2023; Zelikman et al., 2022). Yang et al. (2024) investigated self-distillation as a way to alleviate distribution shifts, improving model quality during SFT while improving generalization across tasks. Our self-data distillation method builds on these techniques by leveraging the original unpruned model to generate a distilled dataset for fine-tuning the pruned model. This enhances semantic alignment and mitigates the quality degradation seen after pruning. Furthermore, while our approach can be combined with KD methods to enhance generalization and recover quality while lowering computational costs, we leave the exploration of such combinations for future work.
@@
One of the major challenges of pruning and distillation techniques in LLMs is catastrophic forgetting, where a model loses its previously learned capabilities during fine-tuning (Kotha et al., 2024; Korbak et al., 2022). Regularization techniques such as Elastic Weight Consolidation (Kirkpatrick et al., 2017) aim to alleviate this by controlling parameter updates, but are task-dependent and require careful tuning (Huang et al., 2021). Architecture-based methods, which allocate separate parameters for each task (Razdaibiedina et al., 2023), preserve task-specific knowledge but add complexity and overhead, reducing the overall efficiency of model compression. Replay-based techniques (Ostapenko et al., 2022; Rolnick et al., 2019; Sun et al., 2019) store data subsets from previous tasks for rehearsal, either through direct storage or synthesis via generative models. However, these methods demand substantial memory to store large datasets and are often impractical due to privacy concerns or lack of access to past data. Our self-data distilled fine-tuning approach avoids these challenges by aligning the fine-tuning dataset with the original model’s learned distribution, preserving knowledge across tasks without requiring new parameters or architectural changes. This method offers a robust solution for mitigating catastrophic forgetting while maintaining model quality after pruning.